OpenStudy (anonymous):

Tiffany kicks a soccer ball off the ground and in the air with an initial velocity of 28 feet per second. Using the formula H(t) = −16t2 + vt + s, what is the maximum height the soccer ball reaches? 15.6 feet 14.1 feet 13.5 feet 12.3 feet

OpenStudy (anonymous):

use v^2 = u^2 + 2gs 0 = 28^2 -64s s = 28^2 / 64 = 12.3 ft